# Random close packing fraction of bidisperse discs: Theoretical derivation and exact bounds

### v1

@@ Abstract
A long-standing problem has been a theoretical prediction of the densest packing fraction of random packings, $\Phi_{RCP}$, of same-size discs in $d=2$ and spheres in $3$. However, to minimize order, experiments and numerical simulations often use two-size discs. For practical purposes, then, a predictive theory for the packing fraction, $\Phi_{RCP}$, of the densest such bidisperse packings is more useful. A disorder-guaranteeing theory is formulated here to fill this gap, using an approach that led to an exact solution for monodisperse discs in $d=2$ Bl21 . $\Phi_{RCP}$ depends on the sizes ratio, $D$, and concentrations, $p$, of the disc types and the developed theory enables derivation of exact and rigorous upper and lower bounds on $\Phi_{RCP}(p,D)$, as well as an explicit prediction of it.
@@
While the problem of the monodisperse RCP appears resolved for discs in $d=2$, many simulations and experiments are carried out on packings of two-size, or bidisperse, discs to minimize crystallization ViBo72 ; Bietal86 ; OkOd04 ; Xuetal05 ; Koetal16 ; Br25 . Therefore, predicting $\Phi_{RCP}$ theoretically in such systems is, arguably, more useful. The bidisperse problem is more complex because, in addition to the difficulties of the infinite process parameter space and limiting order, $\Phi_{RCP}$ in such packings is not a number but a function of the disc sizes ratio, $D$, and the concentration of the smaller discs, $p$. Nevertheless, the approach developed for monodisperse packings paves the way to construct this more involved theory. Developing this approach is the main aim here. The following is limited to discs of $\left(1,D_{max}\right)$ in $d=2$, where $D_{max}=\sqrt{3}/(2-\sqrt{3})$ ensures no ‘rattlers’ within enclosures of three large discs.
@@
The COD: Without loss of generality, the diameter of the smaller discs is taken to be unity and their concentration $p$. Any theory, aiming to predict $\Phi_{RCP}$, faces the problem of sensitivity to the packing process, which exists in an infinite parameter space. This difficulty is alleviated by using the COD MaBl14 ; Waetal20 , defined as follows. Connecting the centres of discs in contact generates a graph, whose nodes are the disc centers and edges are the lines connecting them. The smallest voids enclosed by the edges are the ’cells’, the number of edges surrounding a cell is its order, and its distribution is the COD,
@@
The first advantage of the COD is that it correlates directly with the density - increasing the fraction of low-order cells increases the mean number of contacts per disc, the density, and the packing fraction. Secondly, as any packing process generates a COD, it effectively parameterizes all possible packings. This circumvents the infinite parameter space problem and predictions based on the COD hold for all processes. Thirdly, it can be used to determine directly the packing fraction of any packing of $N$ discs and $N_{c}$ cells, as follows.
@@
Exact upper bound on $\Phi_{RCP}$: The densest possible packings contain only $3$-cells. While ordered such packings can be generated Fe25 , it is unclear whether disordered ones are topologically possible. Nevertheless, such hypothetical packings provide a rigorous upper bound for $\Phi_{RCP}(p,D)$.
@@
For $D\neq 1$, $u\neq p$ because adjacent $3$-cells are correlated through the sharing of two discs. The exact relation $p(u)$ is derived below. As $N\to\infty$, the total area of this packing is
@@
$\Phi_{3}$ also provides the actual packing fraction of ordered $3$-cell bidisperse disc packing, each of which corresponding to a specific combination of $p$, $D$, and $Q_{3}$ Fe25 .
Figure 2: The packing fraction, $\Phi_{3}(p,D)$, for $1.0\leq D\leq 6.0<D_{max}=\sqrt{3}/(2-\sqrt{3})$. The blue line follows the highest packing fraction for each mixture (add ’max’ in its legend). $D=1.o$ recovers the fully trigonal crystal, $\Phi_{3}=\pi/\left(2\sqrt{3}\right)$.
A disorder criterion: Order often appears in the form of large clusters of trigonal lattices of identical $3$-cells, i.e., of clusters comprising only either configuration $a$ or $d$ in Fig. 1. The occurrence probability of such clusters increases with $Q_{3}$ and, close to $p=0$ or $1$, such packings cannot be considered disordered. To identify the allowed range of $p$ within which disorder is assured, a criterion, developed for monodisperse packings, is extended next. It states that, if the probability that a $3$-cell has, on average, only one identical neighboring cell, then the occurrence probability of clusters of $L$ such cells, $P(L)$, decays exponentially with $L$, with $P(L>5)<1.73\times 10^{-4}$ Bl21 . The occurrence probabilities of configurations $a$ and $d$ are, respectively, $u^{3}Q_{3}$ and $v^{3}Q_{3}$ and the probability of a cell of either configuration having more than one identical neighbor is
@@
The densest packing fraction: Anticipating that disordered packings of only $3$-cells may be topologically impossible, the ideal packing should comprise only $3$- and $4$-cells, whose packing fractions are
@@
Within a central region of $p$, $\Phi_{RCP}$ coincides with the upper bound for all $D$. This is because $Q_{3max}=1$ there and those packings are disordered even if they could comprise, hypothetically, only $3$-cells. This issue is discussed in more detail below. Significantly, an inspection of the plots reveals that, for every size ratio, the maximal possible value of $\Phi_{RCP}$ coincides with the upper bound. An observation that should aid planning compositions of bidisperse packings.
Exact lower bound on $\Phi_{RCP}$: Focusing on packings of only $3$- and $4$-cells, it should be noted that the packing are disordered for all $1\leq D<D_{max}$ and any value of $p$ when $Q_{3max}=0.562236$.... As $\Phi$ is a monotonically increasing function of $Q_{3max}$, substituting this value in relation (12) provides an exact and rigorous lower bound on $\Phi_{RCP}$, irrespective of $D$ (even when $D\to 1$),
@@
To conclude, a theory, based on the cell order distribution (COD), has been developed to solve analytically the long-standing problem of random close packing of bidisperse discs in $d=2$, for size ratios that exclude rattlers in $3$-cells. A disorder-assuring criterion has been derived, which is an extension of the one developed for monodisperse disc packings Bl21 . This criterion allows to determine, for any disc size ratio, $D$, the range of small disc concentration, $p$, within which the packing is disordered. This information is useful for avoiding crystallization when designing experiments and simulations of disordered disc systems. In particular, it was shown that a common choice of bidisperse systems whose two disc types occupy the same area, increases the risk of crystallisation for $D\lesssim 3$. Exact $p$- and $D$-dependent upper and lower bounds on $\Phi_{RCP}$ have been derived, the former corresponding to a disordered packing of only $3$-cells and the latter corresponding to the highest fraction of $3$-cells at which disorder is assured for all $D$. Then the exact value of the random close packing fraction, $\Phi_{RCP}$, was derived as a function of $p$ and $D$. It has been found that, for every size ratio, the highest possible packing fraction always coincides with the upper bound. The packing fraction, at which this occurs, $p_{3max}$, can serve as a guide to future physical and numerical experiments in such systems. The use of the COD obviates the sensitivity to the packing process, which introduces an infinite parameter space that is impossible to explore fully by trial and error. It follows that the results obtained here are valid in general, irrespective of the packing protocol.
In setting the disorder criterion, only trigonal order has been considered. While could be argued that large clusters of some $4$-cells may also represent order, being deformed square lattices, it is shown in the supplemental material SM that a disorder criterion, similar to (11), can also be derived for $Q_{4}$ and that the ideal COD determined above always satisfies it.
@@
Finally, the method developed here is useful beyond predicting the densest disordered bidisperse disc packings. As has been noted in passing, the method yields the packing fractions of ordered packings that comprise of only $3$-cells Fe25 . Moreover, it can be used to analyze general multidisperse disc packings of any COD. Such analyses require determining the statistics of cell orders higher than $4$ and, while this could be demanding to achieve analytically or in close form, such calculations can be readily done numerically.

### v2

@@ Abstract
A long-standing problem has been a theoretical prediction of the densest packing fraction of random packings, $\Phi_{RCP}$, of same-size discs in $d=2$ and spheres in 33. However, to minimize order, experiments and numerical simulations often use two-size discs. For practical purposes, then, a predictive theory for the packing fraction, $\Phi_{RCP}$, of the densest such bidisperse packings is more useful. A disorder-guaranteeing theory is formulated here to fill this gap, using an approach that led to an exact solution for monodisperse discs in $d=2$ Bl21 . $\Phi_{RCP}$ depends on the sizes ratio, $D$, and concentrations, $p$, of the disc types and the developed theory enables derivation of exact and rigorous upper and lower bounds on $\Phi_{RCP}(p,D)$, as well as an explicit prediction of it.
@@
While the problem of the monodisperse RCP appears resolved for discs in $d=2$, many simulations and experiments are carried out on packings of two-size, or bidisperse, discs to minimize crystallization ViBo72 ; Bietal86 ; OkOd04 ; Xuetal05 ; Koetal16 ; Br25 . Therefore, predicting $\Phi_{RCP}$ theoretically in such systems is, arguably, more useful. The bidisperse problem is more complex because, in addition to the difficulties of the infinite process parameter space and limiting order, $\Phi_{RCP}$ in such packings is not a number but a function of the disc sizes ratio, $D$, and the concentration of the smaller discs, pp. Nevertheless, the approach developed for monodisperse packings paves the way to construct this more involved theory. Developing this approach is the main aim here. The following is limited to discs of $\left(1,D_{max}\right)$ in $d=2$, where $D_{max}=\sqrt{3}/(2-\sqrt{3})$ ensures no ‘rattlers’ within enclosures of three large discs.
@@
The COD: Without loss of generality, the diameter of the smaller discs is taken to be unity and their concentration pp. Any theory, aiming to predict $\Phi_{RCP}$, faces the problem of sensitivity to the packing process, which exists in an infinite parameter space. This difficulty is alleviated by using the COD MaBl14 ; Waetal20 , defined as follows. Connecting the centres of discs in contact generates a graph, whose nodes are the disc centers and edges are the lines connecting them. The smallest voids enclosed by the edges are the ’cells’, the number of edges surrounding a cell is its order, and its distribution is the COD,
@@
The first advantage of the COD is that it correlates directly with the density - increasing the fraction of low-order cells increases the mean number of contacts per disc, the density, and the packing fraction. Secondly, as any packing process generates a COD, it effectively parameterizes all possible packings. This circumvents the infinite parameter space problem and predictions based on the COD hold for all processes. Thirdly, it can be used to determine directly the packing fraction of any packing of NN discs and $N_{c}$ cells, as follows.
@@
Exact upper bound on $\Phi_{RCP}$: The densest possible packings contain only 33-cells. While ordered such packings can be generated Fe25 , it is unclear whether disordered ones are topologically possible. Nevertheless, such hypothetical packings provide a rigorous upper bound for $\Phi_{RCP}(p,D)$.
@@
For $D\neq 1$, $u\neq p$ because adjacent 33-cells are correlated through the sharing of two discs. The exact relation $p(u)$ is derived below. As $N\to\infty$, the total area of this packing is
@@
$\Phi_{3}$ also provides the actual packing fraction of ordered 33-cell bidisperse disc packing, each of which corresponding to a specific combination of $p$, $D$, and $Q_{3}$ Fe25 .
Figure 2: The packing fraction, $\Phi_{3}(p,D)$, for $1.0\leq D\leq 6.0<D_{max}=\sqrt{3}/(2-\sqrt{3})$. The blue line follows the highest packing fraction for each mixture. $D=1.o$ recovers the fully trigonal crystal, $\Phi_{3}=\pi/\left(2\sqrt{3}\right)$.
A disorder criterion: Order often appears in the form of large clusters of trigonal lattices of identical 33-cells, i.e., of clusters comprising only either configuration $a$ or $d$ in Fig. 1. The occurrence probability of such clusters increases with $Q_{3}$ and, close to $p=0$ or 11, such packings cannot be considered disordered. To identify the allowed range of $p$ within which disorder is assured, a criterion, developed for monodisperse packings, is extended next. It states that, if the probability that a 33-cell has, on average, only one identical neighboring cell, then the occurrence probability of clusters of $L$ such cells, $P(L)$, decays exponentially with $L$, with $P(L>5)<1.73\times 10^{-4}$ Bl21 . The occurrence probabilities of configurations $a$ and $d$ are, respectively, $u^{3}Q_{3}$ and $v^{3}Q_{3}$ and the probability of a cell of either configuration having more than one identical neighbor is
@@
The densest packing fraction: Anticipating that disordered packings of only 33-cells may be topologically impossible, the ideal packing should comprise only 33- and 44-cells, whose packing fractions are
@@
Within a central region of $p$, $\Phi_{RCP}$ coincides with the upper bound for all $D$. This is because $Q_{3max}=1$ there and those packings are disordered even if they could comprise, hypothetically, only 33-cells. This issue is discussed in more detail below. Significantly, an inspection of the plots reveals that, for every size ratio, the maximal possible value of $\Phi_{RCP}$ coincides with the upper bound. An observation that should aid planning compositions of bidisperse packings.
Exact lower bound on $\Phi_{RCP}$: Focusing on packings of only 33- and 44-cells, it should be noted that the packing are disordered for all $1\leq D<D_{max}$ and any value of $p$ when $Q_{3max}=0.562236$.... As $\Phi$ is a monotonically increasing function of $Q_{3max}$, substituting this value in relation (12) provides an exact and rigorous lower bound on $\Phi_{RCP}$, irrespective of $D$ (even when $D\to 1$),
@@
To conclude, a theory, based on the cell order distribution (COD), has been developed to solve analytically the long-standing problem of random close packing of bidisperse discs in $d=2$, for size ratios that exclude rattlers in 33-cells. A disorder-assuring criterion has been derived, which is an extension of the one developed for monodisperse disc packings Bl21 . This criterion allows to determine, for any disc size ratio, $D$, the range of small disc concentration, $p$, within which the packing is disordered. This information is useful for avoiding crystallization when designing experiments and simulations of disordered disc systems. In particular, it was shown that a common choice of bidisperse systems whose two disc types occupy the same area, increases the risk of crystallisation for $D\lesssim 3$. Exact $p$- and $D$-dependent upper and lower bounds on $\Phi_{RCP}$ have been derived, the former corresponding to a disordered packing of only 33-cells and the latter corresponding to the highest fraction of 33-cells at which disorder is assured for all $D$. Then the exact value of the random close packing fraction, $\Phi_{RCP}$, was derived as a function of $p$ and $D$. It has been found that, for every size ratio, the highest possible packing fraction always coincides with the upper bound. The packing fraction, at which this occurs, $p_{3max}$, can serve as a guide to future physical and numerical experiments in such systems. The use of the COD obviates the sensitivity to the packing process, which introduces an infinite parameter space that is impossible to explore fully by trial and error. It follows that the results obtained here are valid in general, irrespective of the packing protocol.
In setting the disorder criterion, only trigonal order has been considered. While could be argued that large clusters of some 44-cells may also represent order, being deformed square lattices, it is shown in the supplemental material SM that a disorder criterion, similar to (11), can also be derived for $Q_{4}$ and that the ideal COD determined above always satisfies it.
@@
Finally, the method developed here is useful beyond predicting the densest disordered bidisperse disc packings. As has been noted in passing, the method yields the packing fractions of ordered packings that comprise of only 33-cells Fe25 . Moreover, it can be used to analyze general multidisperse disc packings of any COD. Such analyses require determining the statistics of cell orders higher than 44 and, while this could be demanding to achieve analytically or in close form, such calculations can be readily done numerically.